Solve the following system by using either addition or substitution. If a unique solution
does not exist, state whether the system is dependent or inconsistent.
10x + 2y = 7
y = -5x + 3

Answer by venugopalramana(3286)

Solve the following system by using either addition or substitution. If a unique solution
does not exist, state whether the system is dependent or inconsistent.
10x + 2y = 7.............................I
y = -5x + 3...OR.......5X+Y=3......................II
2*EQN.II GIVES
10X+2Y=6....WHERE AS EQN.I SAYS THAT 10X+2Y=7...SO THESE ARE INCONSISTENT EQNS.
THERE IS NO POSSIBLE SOLUTION AS IF WE TRY TO SATISFY EQN.I..THEN EQN.II WILL
NOT BE SATISFIED AND VICEVERSA